OpenStudy (anonymous):

Find the percent of change in these numbers. 60 to 75

OpenStudy (anonymous):

\[60=100percent\] \[75=?percent\] Cross multiply to find the question mark.

OpenStudy (anonymous):

25% (Or 125%?). The difference between 60 and 75 is 15, and 15 is 25% of 60.

jimthompson5910 (jim_thompson5910):

Similar Problem (numbers have been changed) Find the percent of change in these numbers: 80 to 120 Percent Increase = [(New - Old)/Old] * 100 Percent Increase = [(120-80)/80] * 100 Percent Increase = [40/80] * 100 Percent Increase = 0.5 * 100 Percent Increase = 50%